OpenStudy (erniellyd28):

8/13-3 to the second power/4+2/3

OpenStudy (anonymous):

9/50

OpenStudy (erniellyd28):

hi how are u thats the answer?

OpenStudy (anonymous):

im good. and if I interpreted the question right, that should be it.

OpenStudy (erniellyd28):

cuz in the book im getting different answers

OpenStudy (anonymous):

could you insert it in the equation button below to see if i interpreted it right?

OpenStudy (erniellyd28):

8 3 2 - - -- + -- 13 4 3

OpenStudy (erniellyd28):

but next to the 3 that reads 3/4 is second power

OpenStudy (anonymous):

so, \[\frac{ 8 }{ 13 } ^\frac{ 3 }{ 4 } + \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

then there is no answer. there are no common denominators..

OpenStudy (erniellyd28):

no sir the problem is written the way i wrote it but next to the 3 where it says 3/4 there is second power

OpenStudy (anonymous):

OH! ok.

OpenStudy (anonymous):

\[\frac{ 8 }{ 13 } - \frac{ 3 }{ 4 }^2 + \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

then it is 449/624

OpenStudy (erniellyd28):

thank u so very much for all your help u have a good day :)